OpenStudy (anonymous):

4+|r+2|=7 Please help and show the steps to answering the problem? Thank you!

OpenStudy (gorv):

subtract 4fromboth side

OpenStudy (anonymous):

|r + 2| = 3

OpenStudy (gorv):

|r+2|=7

OpenStudy (gorv):

now to remove modulus it will give both positive n negative value

OpenStudy (gorv):

\[r+2=\pm7\]

OpenStudy (gorv):

r+2=7 r=7-2=5 r+2=-7 r=-7-2=-9

OpenStudy (anonymous):

its 3 not 7 check again